OpenStudy (anonymous):

researchers studying fish populations in sunset lake caught marked and then released 232 bass. later a sample of 329 bass were caught and examined. of these 16 were found to be marked. use the proportion below to estimate the total bass population in sunset lake. marked bass in sample/total bass in sample = marked bass in lake/total bass in lake

OpenStudy (anonymous):

help please!!!

OpenStudy (anonymous):

16/329 = 232/x 16x = 329*232 16x = 76328 x = 4770.5,

OpenStudy (anonymous):

thanks!!!!!!!!!!!

OpenStudy (anonymous):

Sure Thing (: